# Long-time asymptotics of (1,3)-sign solitary waves for the damped nonlinear Klein-Gordon equation

### v1

@@ Abstract.
We consider the damped nonlinear Klein-Gordon equation:
@@
where $\alpha>0$, $2\leq d\leq 5$ and energy sub-critical exponents $p>2$. In this paper, we prove that any solution which is asymptotic to a superposition of four solitons with exactly one soliton of opposite sign evolves so that the three like-signed solitons spread out in an equilateral-triangle configuration centered at the oppositely signed soliton.
@@
Nonlinear Klein-Gordon equation, Solitons, Soliton resolution, Multi-solitons, Long-time asymptotics
@@ 1.1. Setting of the problem
We consider the following damped nonlinear Klein-Gordon equation:
$\left\{\begin{aligned} &\partial_{t}^{2}u-\Delta u+2\alpha\partial_{t}u+u-f(u)=0,&(t,x)\in\mathbb{R}\times\mathbb{R}^{d},\\ &\left(u(0,x),{\partial}_{t}u(0,x))=(u_{0}(x),u_{1}(x)\right)\in\mathcal{H},\end{aligned}\right$. (DNKG)
@@
It follows from [3] that the Cauchy problem for (DNKG) is locally well-posed in the energy space $\mathcal{H}$. Moreover defining the energy of a solution $\vec{u}=(u,{\partial}_{t}u)$ by
@@
1.2. Long-time asymptotics of multi solitary waves
In this paper, we are interested in the dynamics of multi-solitary waves related to the ground state $Q$, which is the unique positive, radial $H^{1}(\mathbb{R}^{d})$ solution of
@@ 1.2. Long-time asymptotics of multi solitary waves
The ground state generates the stationary solution $(Q,0)$. By (1.2), the function (Q,0)(-Q,0) as well as any translate $(Q(\cdot+y),0)$ are also solutions of (DNKG). There have been intensive studies on global behavior of general (large) solutions for nonlinear dispersive equations, where the guiding principle is the soliton resolution conjecture, which claims that generic global solutions are asymptotic to superposition of solitons for large time. For (DNKG), Feireisl [17] proved the soliton resolution along a time sequence: for any global bounded solution $\vec{u}$ of (DNKG), there is a time sequence $t_{n}\to\infty$ and a set of stationary solutions $\{\varphi_{k}\}_{k=1}^{K}$ ($K\geq 0$), and sequences $\{c_{k,n}\}_{k=1,2,\ldots,K,n\in\mathbb{N}}$ such that
@@
In particular, in the $d=1$ case, since the stationary solutions of (DNKG) consist only of translates of $\pm Q$, Côte, Martel, and Yuan [9] proved that if $\vec{u}$ is a global solution of (DNKG), there exist a sign $\sigma=\pm 1$, $K\geq 0$, and functions $z_{k}:[0,\infty)\to\mathbb{R}^{d}$ such that
$\lim_{t\to\infty}\left(\|\vec{u}(t)-\sigma\sum_{k=1}^{K}(-1)^{k}Q(\cdot-z_{k}(t))\|_{H^{1}}+\|{\partial}_{t}u(t)\|_{L^{2}}\right)$ $=0,$
$\lim_{t\to\infty}\left|z_{k+1}(t)-z_{k}(t)\right|$ $=\infty.$
@@
Prior to the present result, Côte, Martel, Yuan, and Zhao [10] proved in arbitrary space dimension that any two-soliton superposition built from $Q$ must consist of solitons of opposite signs, and that the solitons move so that their centers evolve along a fixed straight line. This result, in particular, makes the interaction forces acting on the solitons explicit, and suggests that like-signed solitons attract whereas oppositely signed solitons repel. Using this force law, [9] showed that in one space dimension, in a multi-soliton superposition the signs of neighboring solitons must be opposite.
@@
On the other hand, when the space dimension is at least two, two types of difficulties arise. The first is that, in dimensions $d\geq 2$, there exist many different kinds of stationary solutions (here “different” means inequivalent even after allowing for spatial translations and sign changes). The second is that the way the configuration spreads in space becomes genuinely directional. When considering a superposition of copies of $Q$, each soliton feels an attractive or repulsive interaction from other nearby solitons. In [9], thanks to the restriction to one space dimension, the centers of the solitons in such a superposition lie on a single line, so it suffices to examine the interactions with adjacent solitons. By contrast, in dimensions $d\geq 2$, three or more solitons can be simultaneously close, and one must analyze not only the distances but also the relative directions. Despite these difficulties, there has been notable recent progress on multi-soliton solutions that can be described as a superposition of copies of $Q$. In arbitrary space dimension, Côte, Martel, Yuan, and Zhao [10] clarified that a two-soliton superposition experiences essentially no “rotational” (i.e., tangential) interaction, and that the inter-soliton force is attractive for like signs and repulsive for opposite signs. Subsequently, the author [21] showed that solutions corresponding to a superposition of three solitons are also collinear. Indeed, unless the three solitons are collinear, the oppositely signed soliton is driven outward by the repulsion from the other two, while the attraction between the like-signed solitons decreases the effective energy of the soliton component; hence the configuration cannot sustain the energy of a three-soliton state. Moreover, as a way of constructing multi-soliton solutions, Côte and Du [7] imposed symmetry conditions on the solution and constructed solutions whose soliton centers form highly symmetric configurations, such as an equilateral triangle or a regular tetrahedron.
@@ 1.3. Main result
First, we define a solution that behaves as the superposition of ground states.
@@
Since (DNKG) is invariant under the sign change $u\mapsto-u$, any property established for $(m,n)$-sign solitary waves immediately yields the corresponding property for $(n,m)$-sign solitary waves. It was also proved in [7] that neither $(m,0)$-sign solitary waves nor $(0,m)$-sign solitary waves exist.
@@
In this paper, we are interested in $(1,3)$-sign solitary waves. By imposing suitable symmetries, Côte and Du [7] constructed $(1,3)$-sign solitary waves for which the center of the +Q+Q-soliton converges to a point, while the centers of the remaining three solitons spread out so as to form an expanding equilateral triangle. In the present work, we prove that this geometric behavior does not rely on any symmetry assumption: if a solution is a $(1,3)$-sign solitary waves, then the center of the +Q+Q-soliton converges, and the other three soliton centers diverge in an equilateral-triangle configuration with the +Q+Q-soliton as the center of similarity. Before stating our main theorem, we define spreading out in an equilateral-triangle configuration by introducing the following terminology.
@@
If $\omega_{1},\omega_{2},\omega_{3}\in S^{d-1}$ form an equilateral triple, then they lie in a common two-dimensional subspace of $\mathbb{R}^{d}$ (indeed, $\omega_{3}=-(\omega_{1}+\omega_{2})$). Moreover, the triangle with vertices $\omega_{1},\omega_{2},\omega_{3}$ is equilateral.
@@
If a solution $\vec{u}$ of (DNKG) is a $(1,3)$-sign solitary waves, there exist $z_{\infty}\in\mathbb{R}^{d}$, an equilateral triple $\omega_{1},\omega_{2},\omega_{3}\in S^{d-1}$, and a function $z_{k}:[0,\infty)\to\mathbb{R}^{d}$ for $k=1,2,3$ such that for all $t>2$,
@@
If $\vec{u}$ is a $(3,1)$-sign solitary-wave solution, then $\vec{u}$ is a $(1,3)$-sign solitary-wave solution by (1.2). Therefore, by Theorem 1.1, the long-time dynamics of the soliton centers for $(3,1)$-sign solitary waves are the same as those for $(1,3)$-sign solitary waves.
@@
To completely characterize the long-time dynamics of four solitary waves solutions, it is also necessary to understand the long-time behavior of $(2,2)$-sign solitary waves. In contrast with the $(1,3)$-sign case, the $(2,2)$-sign configuration has no distinguished (reference) soliton—there is no unique soliton whose sign differs from all the others—and this makes the asymptotic analysis more delicate.
@@ 1.4. Difficulties and Ideas for the main result
Determining the long-time behavior of a four-soliton configuration is more difficult than in the case of a superposition of three or fewer solitons. Indeed, since each soliton is acted upon by forces generated by the other three solitons, one cannot a priori rule out the possibility that the soliton centers exhibit complicated long-time motions. For example, in the three-soliton case, [21] shows that it suffices to prove first that the distances between solitons of opposite signs become close to one another; under such a situation, if the three centers are not collinear, then the oppositely signed soliton is pushed outward by the repulsive forces exerted by the other two solitons, and the angle of the corresponding isosceles triangle closes. By contrast, in the $(1,3)$-sign solitary waves setting considered here, the forces act from three different directions, so that even the force exerted on the oppositely signed soliton by two of the solitons may be canceled by the force coming from the remaining soliton. Because the inter-soliton forces are so delicate, it is conceivable a priori that, for $(1,3)$-sign solitary waves, the three like-signed solitons might spread out in a non-equilateral configuration—for instance, forming a triangle with angles $50^{\circ}$, $60^{\circ}$, and $70^{\circ}$—rather than an equilateral one. In this paper, to study the motion of the soliton centers, we exploit the fact that the mutual distances between solitons evolve in a way that is coupled to the geometry of the configuration, in particular to the angles formed by the centers. More precisely, taking the oppositely signed soliton as a reference, we consider the three directions along which the other solitons separate, and introduce the associated unit direction vectors; the angles between these directions are encoded by the inner products of these unit vectors. We then derive an effective system of ordinary differential equations for these angular variables and reduce it to a closed model ODE that can be written solely in terms of the three pairwise inner products. The resulting dynamical system can be analyzed by an elementary study of its trajectories. A key input behind this reduction is a separation of time scales: the angular variables evolve much more slowly than the radial separations, so that, to leading order, one may treat the distances as evolving as if the angles were frozen.
@@ 1.5. Previous results
First, we introduce results related to the soliton resolution conjecture. The soliton resolution conjecture has been studied for the energy-critical wave equation. We refer to [15, 27], in which the conjecture was proven for all radial solutions in space dimension $d\geq 3$. Historically, we refer to [6, 13, 15, 16], where the conjecture has been proved in space dimension $d=3,5,7,\ldots$ or $d=4,6$ using the method of channels. In recent years, Jendrej and Lawrie [27] proved the conjecture in space dimension $d\geq 4$ using a novel argument based on the analysis of collision intervals. Similar results regarding the conjecture have been obtained for cases with a potential [29, 35], with a damping term [19], as well as for the wave maps [14, 25], for the heat equation [2], all under some rotational symmetry.
On the other hand, the conjecture remains widely open for the nonlinear Klein-Gordon equation, namely the $\alpha=0$ case. Nakanishi and Schlag [39] proved the conjecture as long as the energy of the solution is not much larger than the ground state energy in space dimension $d=3$. For the damped nonlinear Klein-Gordon equation case, the energy decay (1.1) makes the analysis simpler than the undamped case. Historically, Keller [30] constructed stable and unstable manifolds in a neighborhood of any stationary solution of (DNKG), and Feireisl [17] proved the conjecture along a time sequence. Burq, Raugel and Schlag [3] proved the conjecture in the full limit $t\to\infty$ for all radial solutions. In particular, Côte, Martel and Yuan [9] proved the conjecture for $d=1$ without size restriction and radial assumption for a non-integrable equation. In a result without any size or symmetry restriction, Kim and Kwon [31] proved the conjecture for Calogero-Moser derivative nonlinear Schödinger equation.
On the other hand, in the case of general dimensions $d\geq 2$, it remains unclear how solitary waves are arranged and how they move. This is because the ODE system for the centers of the solitary waves becomes more complex in general dimensions.
@@
In addition, several studies have been conducted on multi-solitary waves for the nonlinear dispersive equations. For the nonlinear Klein-Gordon equation, Côte and Muñoz [11] constructed multi-solitary waves based on the ground state. Furthermore, Côte and Martel [8] extended the work of [11] by constructing multi-solitary waves based on the stationary solution of the nonlinear Klein-Gordon equation. Notably, multi-solitary waves usually behave asymptotically as the solitary waves moving at distinct constant speeds. More recently, Aryan [1] has constructed a 2-solitary wave where the distance between each solitary wave is $2(1+O(1))\log{t}$ as $t\to\infty$. There are several results regarding multi-solitary waves with logarithmic distance; see e.g. [18, 36, 37, 38]. Notably, for the damped nonlinear Klein-Gordon equation, multi-solitary waves of (DNKG) cannot move at a constant speed by the damping $\alpha$. [23, 24, 25, 26, 27, 28] are studies on two-bubble solutions.

### v2

@@ Abstract.
We consider the damped nonlinear Klein-Gordon equation
@@
where $\alpha>0$, $2\leq d\leq 5$ and energy sub-critical exponents $p>2$. In this paper, we prove that any solution that is asymptotic to a superposition of four solitons with exactly one soliton of opposite sign evolves so that the three like-signed solitons spread out in an equilateral-triangle configuration centered at the oppositely signed soliton.
@@
Nonlinear Klein-Gordon equation, Solitons, Soliton resolution, Multi-solitons, Long-time asymptotics, Regular polygon, Rigidity
@@ 1.1. Setting of the problem
We consider the following damped nonlinear Klein-Gordon equation
$\left\{\begin{aligned} &\partial_{t}^{2}u-\Delta u+2\alpha\partial_{t}u+u-f(u)=0,&(t,x)\in\mathbb{R}\times\mathbb{R}^{d},\\ &\left(u(0,x),{\partial}_{t}u(0,x)\right)=\left(u_{0}(x),u_{1}(x)\right)\in\mathcal{H},\end{aligned}\right$. (DNKG)
@@
It follows from [4] that the Cauchy problem for (DNKG) is locally well-posed in the energy space $\mathcal{H}$. Moreover, defining the energy of a solution $\vec{u}=(u,{\partial}_{t}u)$ by
@@
1.2. Long-time asymptotics of multi-soliton solutions
In this paper, we are interested in the dynamics of multi-soliton solutions related to the ground state $Q$, which is the unique positive, radial $H^{1}(\mathbb{R}^{d})$ solution of
@@ 1.2. Long-time asymptotics of multi-soliton solutions
(See [31]). The ground state generates the stationary solution $(Q,0)$. By (1.2), the function (Q,0)(-Q,0) as well as any translate $(Q(\cdot+y),0)$ is also a solution of (DNKG). There have been intensive studies on global behavior of general (large) solutions for nonlinear dispersive equations, where the guiding principle is the soliton resolution conjecture, which claims that generic global solutions are asymptotic to superposition of solitons for large time. For (DNKG), Feireisl [16] proved the soliton resolution along a time sequence: for any global bounded solution $\vec{u}$ of (DNKG), there exist a time sequence $t_{n}\to\infty$ and a set of stationary solutions $\{\varphi_{k}\}_{k=1}^{K}$ ($K\geq 0$), and sequences $\{c_{k,n}\}_{k=1,2,\ldots,K,n\in\mathbb{N}}$ such that
@@
In particular, in the $d=1$ case, since the stationary solutions of (DNKG) consist only of translates of $\pm Q$, Côte, Martel, and Yuan [8] proved that if $\vec{u}$ is a global solution of (DNKG), there exist a sign $\sigma=\pm 1$, $K\geq 0$, and functions $z_{k}:[0,\infty)\to\mathbb{R}$ such that
$\lim_{t\to\infty}\left(\|u(t)-\sigma\sum_{k=1}^{K}(-1)^{k}Q(\cdot-z_{k}(t))\|_{H^{1}}+\|{\partial}_{t}u(t)\|_{L^{2}}\right)$ $=0,$
$\lim_{t\to\infty}\left(z_{k+1}(t)-z_{k}(t)\right)$ $=\infty.$
@@
Prior to the present result, Côte, Martel, Yuan, and Zhao [9] proved in general space dimension that any two-soliton superposition built from $Q$ must consist of solitons of opposite signs, and that the solitons move so that their centers evolve along a fixed straight line. In particular, this result makes the interaction forces acting on the solitons explicit, and suggests that like-signed solitons attract whereas oppositely signed solitons repel. Using this force law, [8] showed that in one space dimension, in a multi-soliton superposition the signs of neighboring solitons must be opposite.
@@
On the other hand, when the space dimension is at least two, two types of difficulties arise. The first is that, in dimensions $d\geq 2$, there exist many different kinds of stationary solutions (here “different” means inequivalent even after allowing for spatial translations and sign changes). The second is that soliton configurations depend on relative directions as well as distances. When considering a superposition of copies of $Q$, each soliton feels an attractive or repulsive interaction from other nearby solitons. In [8], thanks to the restriction to one space dimension, the centers of the solitons in such a superposition lie on a single line, so it suffices to examine the interactions with adjacent solitons. By contrast, in dimensions $d\geq 2$, three or more solitons can be simultaneously close, and one must analyze not only the distances but also the relative directions. Despite these difficulties, there has been notable recent progress on multi-soliton solutions that can be described as superpositions of copies of $Q$. As mentioned above, Côte, Martel, Yuan, and Zhao [9] showed that, in general space dimension, a two-soliton superposition experiences essentially no “rotational” (i.e., tangential) interaction, and that the inter-soliton force is attractive for like signs and repulsive for opposite signs. Subsequently, the author [20] showed that three-soliton superpositions are necessarily collinear. Indeed, it is known that any solution asymptotic to a superposition of three solitons must contain exactly one soliton whose sign is opposite to that of the other two (see [6]). In this case, unless the three solitons are collinear, the oppositely signed soliton is pushed outward by repulsion from the other two; consequently, the attraction between the like-signed solitons dominates the repulsion between oppositely signed solitons, and the configuration cannot sustain a three-soliton state. We also note that, as a constructive approach, Côte and Du [6] imposed symmetry conditions and constructed solutions whose soliton centers form highly symmetric configurations, such as an equilateral triangle or a regular tetrahedron.
@@ 1.3. Main result
First, we define solutions that behave as the superposition of ground states.
@@
Since (DNKG) is invariant under the sign change $u\mapsto-u$, any property established for $(m,n)$-sign solitary waves solutions immediately yields the corresponding property for $(n,m)$-sign solitary waves solutions. It was also proved in [6] that no $(m,0)$-sign or $(0,m)$-sign solitary waves solutions exist.
@@
In this paper, we are interested in $(1,3)$-sign solitary waves solutions. By imposing suitable symmetries, Côte and Du [6] constructed a $(1,3)$-sign solitary waves for which the center of the +Q+Q-soliton converges to a point, while the centers of the remaining three solitons spread out so as to form an expanding equilateral triangle. In the present work, we prove that this geometric behavior does not rely on any symmetry assumption: if a solution is a $(1,3)$-sign solitary waves, then the center of the +Q+Q-soliton converges, and the other three soliton centers diverge in an equilateral-triangle configuration with the +Q+Q-soliton as the center of similarity. Before stating our main theorem, we define spreading out in an equilateral-triangle configuration by introducing the following terminology.
@@
If $\omega_{1},\omega_{2},\omega_{3}\in S^{d-1}$ form an equilateral triple, then they lie in a common two-dimensional subspace of $\mathbb{R}^{d}$ (indeed, $\omega_{3}=-(\omega_{1}+\omega_{2})$). Moreover, $\omega_{1},\omega_{2},\omega_{3}$ satisfy
$|\omega_{1}-\omega_{2}|=|\omega_{2}-\omega_{3}|=|\omega_{3}-\omega_{1}|.$
@@
If a solution $\vec{u}$ of (DNKG) is a $(1,3)$-sign solitary waves, there exist $z_{\infty}\in\mathbb{R}^{d}$, an equilateral triple $\omega_{1},\omega_{2},\omega_{3}\in S^{d-1}$, and functions $z_{k}:[0,\infty)\to\mathbb{R}^{d}$ for $k=1,2,3$ such that for all $t>2$,
@@
If $\vec{u}$ is a $(3,1)$-sign solitary waves, then $\vec{u}$ is a $(1,3)$-sign solitary waves by (1.2). Therefore, by Theorem 1.1, the long-time dynamics of the soliton centers for a $(3,1)$-sign solitary waves are the same as those for a $(1,3)$-sign solitary waves.
@@
To completely characterize the long-time dynamics of four solitary waves solutions, it is also necessary to understand the long-time behavior of $(2,2)$-sign solitary waves solutions. In contrast to the $(1,3)$-sign case, the $(2,2)$-sign configuration has no distinguished (reference) soliton—there is no unique soliton whose sign differs from all the others—and this makes the asymptotic analysis more delicate.
@@ 1.4. Difficulties and Ideas for the Main Result
Determining the long-time behavior of a four-soliton configuration is more difficult than in the case of a superposition of three or fewer solitons. Indeed, since each soliton is acted on by forces generated by the other three solitons, one cannot rule out a priori the possibility that the soliton centers exhibit complicated long-time motions. For instance, in the three-soliton case, [20] shows that it suffices to first prove that the two distances from the oppositely signed soliton to the other two become comparable. In that regime, if the three centers are not collinear, then the oppositely signed soliton is pushed outward by the repulsive forces exerted by the other two, and the apex angle of the corresponding isosceles triangle decreases. By contrast, in the $(1,3)$-sign solitary waves setting considered here, the forces act from three different directions, so that even the force exerted on the oppositely signed soliton by two of the solitons may be canceled by the force coming from the remaining soliton. Because the inter-soliton forces are so delicate, it is conceivable a priori that, for $(1,3)$-sign solitary waves, the three like-signed solitons might spread out in a non-equilateral configuration—for instance, forming a triangle with angles $50^{\circ}$, $60^{\circ}$, and $70^{\circ}$—rather than an equilateral one. In this paper, to study the motion of the soliton centers, we exploit the fact that the mutual distances between solitons evolve in a way that is coupled to the geometry of the configuration, in particular to the angles formed by the centers. More precisely, taking the oppositely signed soliton as a reference, we consider the three directions along which the other solitons separate, and introduce the associated unit direction vectors; the angles between these directions are encoded by the inner products of these unit vectors. We then derive an effective system of ordinary differential equations for these angular variables and reduce it to a closed ODE system that can be written solely in terms of the three pairwise inner products. The resulting dynamical system can be analyzed by an elementary phase-portrait analysis. A key idea behind this reduction is a separation of time scales: the angular variables evolve much more slowly than the radial distances, so that, to leading order, one may treat the distances as evolving as if the angles were frozen.
@@ 1.5. Previous results
First, we introduce results related to the soliton resolution conjecture. The soliton resolution conjecture has been studied for the energy-critical wave equation. We refer to [14, 27], in which the conjecture was proven for all radial solutions in space dimension $d\geq 3$. Historically, we refer to [5, 12, 14, 15], where the conjecture has been proved in space dimension $d=3,5,7,\ldots$ or $d=4,6$ using the method of channels. In recent years, Jendrej and Lawrie [27] proved the conjecture in space dimension $d\geq 4$ using a novel argument based on the analysis of collision intervals. Similar results regarding the conjecture have been obtained for cases with a potential [28, 32], with a damping term [18], as well as for wave maps [13, 25], for the heat equation [2], all under some rotational symmetry. In recent years, in order to gain a finer understanding of the long-time dynamics, there has been growing interest in the analysis of intermediate (or transient) states that arise before the final soliton-resolution regime as $t\to+\infty$. For energy-critical wave equations, see [38, 39], and for the heat equation, see [3].
On the other hand, the conjecture remains widely open for the nonlinear Klein-Gordon equation, namely the $\alpha=0$ case. Nakanishi and Schlag [37] proved the conjecture as long as the energy of the solution is not much larger than the ground state energy in space dimension $d=3$. For the damped nonlinear Klein-Gordon equation case, the energy decay (1.1) makes the analysis simpler than the undamped case. Historically, Keller [29] constructed stable and unstable manifolds in a neighborhood of any stationary solution of (DNKG), and Feireisl [16] proved the conjecture along a time sequence. Burq, Raugel and Schlag [4] proved the conjecture in the full limit $t\to\infty$ for all radial solutions. In particular, Côte, Martel and Yuan [8] proved the conjecture for $d=1$ without size restriction and radial assumption for a non-integrable equation. In a result without any size or symmetry restriction, Kim and Kwon [30] proved the conjecture for Calogero-Moser derivative nonlinear Schödinger equation. On the other hand, in the case of general dimensions $d\geq 2$, it remains unclear how solitary waves are arranged and how they move. One reason is that, in general dimensions, it is not known whether solutions remain bounded in the energy space $\mathcal{H}$; another is that (DNKG) admits a wide variety of stationary solutions.
@@
In addition, several studies have been conducted on multi-solitary waves for the nonlinear dispersive equations. For the nonlinear Klein-Gordon equation, Côte and Muñoz [10] constructed multi-solitary waves based on the ground state. Furthermore, Côte and Martel [7] extended the work of [10] by constructing multi-solitary waves based on the stationary solution of the nonlinear Klein-Gordon equation. Notably, multi-solitary waves usually behave asymptotically as the solitary waves moving at distinct constant speeds. More recently, Aryan [1] has constructed a 2-solitary wave where the distance between each solitary wave is $2(1+O(1))\log{t}$ as $t\to\infty$. There are several results regarding multi-solitary waves with logarithmic distance; see e.g. [17, 33, 35, 36]. Notably, for the damped nonlinear Klein-Gordon equation, multi-solitary waves of (DNKG) cannot move at a constant speed by the damping $\alpha$. A distinctive feature of the present work is that it provides a complete rigidity description for the long-time dynamics of multi-soliton solutions. Most of the aforementioned results on multi-solitons have focused primarily on construction and existence. Among them, the results in [6] and [34] concern regular polygonal configurations for dispersive equations. In [6], the existence of such a regular polygon configuration is obtained by imposing symmetry assumptions on the solutions. In [34], the authors constructed a multi-bubble (blow-up bubble) solution which is asymptotic to a superposition of concentrating bubbles centered at the vertices of a regular polygon. In both works, symmetry is imposed in order to simplify the interactions between solitons, since without symmetry the interactions become significantly more involved. To the best of my knowledge, the present paper is the first to show, in the context of dispersive equations, that the centers of multi-solitons are forced to form an equilateral triangle.
Remark 1.9.
By a “bubble solution”, we mean a stationary solution (or standing wave) obtained from a fixed stationary profile by applying the scaling symmetry. In this paper, we refer to such solutions as “solitons”. For instance, in the context of soliton resolution for the energy-critical wave equation, the solitons that arise are precisely bubble solutions. Moreover, by a “multi-bubble solution”, we mean a solution that is asymptotic to a superposition of bubbles. We emphasize that the term “multi-bubble solution” is not restricted to the case where all bubbles have the same scale, as in [34]. In many contexts, it also includes superpositions of bubble solutions associated with different scales. Indeed, [22, 23, 24, 25, 26] study two-bubble solutions with different scales.